OpenStudy (anonymous):

The value of 3to the 3rd power +4 to the 2nd power=

OpenStudy (anonymous):

The exponents would come first, so you would start with 3*3*3 and 4*4. When you solve that, add the answers together.

Mehek (mehek14):

\(3^3+4^2\\3^3=27\\4^2=16\\27+16=?\) @ripena24